# Magic Gems: A Polyhedral Framework for Magic Squares Connecting Combinatorics, Geometry, and Linear Algebra

### v1

@@ 4.1.3 Covariance Verification and Geometric Structure
We computed the covariance matrices for all eight $D_{4}$ variants of the Lo Shu square. As predicted by Section 3.6, every variant achieves $\mathrm{Cov}(X,Z)=\mathrm{Cov}(Y,Z)=0$ to machine precision.
Notably, exhaustive enumeration reveals that 760 arrangements of $\{1,\ldots,9\}$ satisfy $\mathrm{Cov}(X,Z)=\mathrm{Cov}(Y,Z)=0$, yet only 8 of these are magic squares. The remaining 752 arrangements fail the diagonal conditions or have unequal row/column sums despite achieving zero row and column covariance. This confirms that the full energy functional $E(S)=0$, not merely the row and column covariances, is required to characterize magic squares.
@@
The $3\times 3$ Magic Gem comprises nine points, of which eight lie on the convex hull while the ninth—corresponding to the center cell with value 5—occupies the interior at the origin. Section 3 presents a comprehensive four-panel analysis of the geometric and statistical structure.
@@
4.2 The 4×4 Case: Even-Order Analysis
@@
A notable difference from the $3\times 3$ case is the grid centering: for even $n$, the cell centers lie at half-integer coordinates ($\pm 0.5,\pm 1.5$), and no cell coincides with the origin. Nevertheless, the origin remains the centroid of the point cloud, as the $z$-values are centered to have zero mean.
Section 3.6 makes no distinction between even and odd orders, predicting $\mathrm{Cov}(X,Z)=\mathrm{Cov}(Y,Z)=0$ for all magic squares. Our computations confirm this prediction for the $4\times 4$ case: the magic square achieves exactly zero covariance to machine precision.
Section 4 presents a comprehensive six-panel analysis of the $4\times 4$ case, including the magic square, Magic Gem, vector representation, construction grid, covariance distributions, and hull volume analysis.
Figure 4: Complete analysis of the $4\times 4$ Magic Gem. (a) The $4\times 4$ magic square constructed via the doubly-even method, with magic constant $M(4)=34$. (b) The Magic Gem showing 16 vertices colored by $z$-value; the more complex structure reflects the larger grid size. (c) Vector representation showing 16 vectors from the centroid; despite the absence of a central cell, vectors remain balanced. (d) Centered grid with cell values at half-integer coordinates. (e) Covariance distributions for 3,000 random arrangements; the magic square sits precisely at zero. (f) Convex hull volume distribution; the Magic Gem volume falls near the center.
4.3 The 5×5 Case: Return to Odd Order
Proceeding to $n=5$ returns us to odd order, where the Siamese construction method applies. The $5\times 5$ case is substantially more complex: there exist approximately 275 million magic squares forming about 34 million equivalence classes.
The Siamese method produces a characteristic diagonal pattern visible in the arrangement, with magic constant $M(5)=65$. The construction maps the 25-cell grid to 25 points in $\mathbb{R}^{3}$, with the central cell again coinciding with the origin.
The $5\times 5$ magic square achieves $\mathrm{Cov}(X,Z)=\mathrm{Cov}(Y,Z)=0$ exactly, as predicted by the theorem. Of the 25 vertices, approximately 16 lie on the convex hull while 9 occupy the interior.
Section 5 presents the complete analysis for the $5\times 5$ case, following the same structure as the $4\times 4$ analysis.
Figure 5: Complete analysis of the $5\times 5$ Magic Gem. (a) The $5\times 5$ magic square constructed via the Siamese method, with magic constant $M(5)=65$. (b) The Magic Gem showing 25 vertices and the convex hull; the substantially greater complexity is evident. (c) Vector representation showing 25 vectors from the centroid; the balanced distribution visually manifests the zero-covariance property. (d) Centered grid with cell values at integer coordinates. (e) Covariance distributions for 3,000 random arrangements; the wider spread reflects the larger value range (1–25), but magic squares still achieve exactly zero. (f) Convex hull volume distribution for random arrangements.
@@ 4.4 Cross-Order Comparison and Scaling Analysis
Section 6 presents a unified view of Magic Gems across all three orders along with a systematic comparison of geometric properties, revealing how the structure evolves with increasing complexity while maintaining the invariant zero-covariance property.
Figure 6: Cross-order comparison and scaling analysis. Top row (a–c): Side-by-side comparison of $3\times 3$, $4\times 4$, and $5\times 5$ Magic Gems. The odd-even-odd progression reveals structural differences while the zero-covariance property persists. Bottom row (d–f): Scaling properties across orders. (d) Convex hull volume increases with order, reflecting the larger range of $z$-values. (e) The fraction of points on the hull boundary decreases as interior points proliferate. (f) The magic constant $M(n)=n(n^{2}+1)/2$ grows quadratically.
Several patterns emerge from this analysis. The convex hull volume grows rapidly with $n$, reflecting the expanding range of $z$-values (4-4 to $4$ for $n=3$, versus 12-12 to $12$ for $n=5$) and the increasing number of vertices. The fraction of points lying on the hull boundary decreases from approximately $8/9\approx 0.89$ for $n=3$ to roughly $0.64$ for $n=5$, as the interior of the hull becomes more populated.
Section 2 summarizes key properties across orders, demonstrating both the growth in complexity and the invariance of the zero-covariance characterization.
@@
Hull vertices 8 $\sim$12 $\sim$16
Interior points 1 $\sim$4 $\sim$9
@@
Table 2: Comparison of Magic Gem properties across orders 3, 4, and 5. Despite differences in parity, construction method, and complexity, all orders exhibit zero covariance as predicted by Section 3.6.
The even-order case ($n=4$) demonstrates that the zero-covariance property does not depend on having a central cell at the origin. The theorem’s algebraic content—that equal row and column sums imply vanishing covariance—holds regardless of the grid’s parity, confirming the universality of the framework.
@@ 4.5 Large-Scale Energy Landscape Analysis
To provide rigorous computational support for the Energy Characterization Theorem and investigate how the energy landscape evolves with order, we conducted extensive sampling of the arrangement space. Leveraging modern multi-core hardware (Mac Studio M5 with 128GB RAM), we analyzed over 460 million arrangements across orders $n=3,4,5$: exhaustive enumeration for $n=3$ (362,880 arrangements), 60 million samples for $n=4$ (from 20.9 trillion total), and 400 million samples for $n=5$ (from approximately $15.5\times 10^{24}$ total).
@@ 4.5.1 Energy Distribution Evolution
Section 7 presents a comprehensive 9-panel analysis of the energy landscape across orders $n=3,4,5$, revealing a dramatic evolution in landscape structure. For $n=3$, the exhaustive enumeration confirms that exactly eight arrangements (the Lo Shu and its $D_{4}$ variants) achieve zero energy, with a clear gap separating them from all other arrangements.
Figure 7: Large-scale energy landscape analysis across orders $n=3,4,5$. Based on exhaustive enumeration ($n=3$: 362,880), and extensive sampling ($n=4$: $6\times 10^{7}$; $n=5$: $4\times 10^{8}$). Top row: Individual distributions (log scale) showing modal energy shifting from 3.12 to 15.12. Middle row: Overlaid comparison, normalized PDFs, and cumulative distributions. Bottom row: Mean energy with error bars, landscape entropy (near-constant at $H\approx 4.4$), and peak shift with quadratic fit ($R^{2}>0.99$). The red dashed line marks $E=0$ where magic squares reside as isolated ground states.
The key statistics from this analysis are summarized in Section 3. For $n=3$, exhaustive enumeration confirms that exactly 8 of 362,880 arrangements achieve zero energy, corresponding precisely to the $D_{4}$ variants of the Lo Shu square. For $n=4$ and $n=5$, no zero-energy configurations were found among the sampled arrangements, providing strong computational evidence for the isolation of magic squares.
@@
Zero-energy fraction $2.2\times 10^{-5}$ $5.8\times 10^{-7}$ $2.8\times 10^{-8}$
Magic squares found 8 0 0
Table 3: Energy landscape statistics from large-scale computational analysis. Based on exhaustive enumeration for $n=3$ and extensive random sampling for $n=4,5$. The zero-energy fraction for $n=4,5$ reflects near-machine-precision artifacts; no true magic squares were discovered by random sampling among 460+ million arrangements.
@@ 4.5.2 Scaling Laws
where $p_{k}$ is the fraction of arrangements falling in energy bin $k$ (using 100 equally-spaced bins covering the observed energy range). Remarkably, the landscape entropy remains nearly constant across orders ($\approx 4.4$$5.0$). This suggests that the relative structure of the landscape is scale-invariant: the difficulty of navigating the landscape by local search methods does not fundamentally change with order, though the absolute energy scales increase substantially.
@@ 4.5.3 Scaling Properties Summary
Section 8 provides a compact four-panel summary of how energy landscape properties scale with magic square order, synthesizing the key quantitative findings. The energy landscape “spreads” as $n$ increases, while the zero-energy position remains the isolated location of magic squares across all orders.
Figure 8: Scaling Properties Across Orders. (A) Peak energy follows quadratic scaling with $R^{2}>0.99$. (B) Mean energy with standard deviation error bars, showing increasing variance. (C) Landscape entropy remains near-constant at $H\approx 4.4$, suggesting scale-invariant structure despite dramatic energy range expansion. (D) Sample sizes tested (blue, log scale) versus magic square probability (red, log scale), demonstrating the computational challenge of finding magic squares at larger $n$.
@@ 4.5.4 Computational Verification Summary
This large-scale analysis provides strong computational support for several key claims:
@@
    Energy Characterization (n=3): Exhaustively verified—exactly 8 of 362,880 arrangements have zero energy, all of which are magic squares. Notably, 760 arrangements satisfy $\mathrm{Cov}(X,Z)=\mathrm{Cov}(Y,Z)=0$, but only the 8 magic squares additionally satisfy the diagonal conditions, confirming that all four covariances must vanish.
@@
    Isolation at Scale: Among 460+ million sampled arrangements across three orders, no false-positive zero-energy configurations were discovered, providing strong evidence that magic squares are the unique ground states.
@@
    Scaling Behavior: The energy landscape exhibits systematic scaling, with the peak energy appearing to grow approximately quadratically with order based on the three data points available ($n=3,4,5$).
@@
These computational results complement the theoretical analysis of Section 3, providing empirical validation of the energy characterization framework and establishing quantitative scaling laws for future investigation.
@@ 5.1 The Meaning of Zero Covariance
The zero-covariance property of magic squares (Section 3.6) admits several complementary interpretations that illuminate their structure.

### v2

@@ 4.1.3 Covariance Verification and Geometric Structure
We computed the covariance matrices for all eight $D_{4}$ variants of the Lo Shu square. As predicted by Proposition 3.6, every variant achieves $\mathrm{Cov}(X,Z)=\mathrm{Cov}(Y,Z)=0$ to machine precision.
Notably, exhaustive enumeration reveals that 760 arrangements of $\{1,\ldots,9\}$ satisfy $\mathrm{Cov}(X,Z)=\mathrm{Cov}(Y,Z)=0$, yet only 8 of these are magic squares. The remaining 752 arrangements fail the diagonal conditions or have unequal row/column sums despite achieving zero row and column covariance. This confirms that all four covariance conditions (including diagonals), not merely the row and column covariances, are required to characterize magic squares.
@@
The $3\times 3$ Magic Gem comprises nine points, of which eight lie on the convex hull while the ninth—corresponding to the center cell with value 5—occupies the interior at the origin. Figure 3 presents a comprehensive four-panel analysis of the geometric and statistical structure.
@@
4.1.4 Extended Statistical Properties
Beyond the zero-covariance result, the Magic Gem coordinates exhibit additional statistical regularities that follow from the magic square structure.
Exact Variance Formulas.
The variances of the Magic Gem coordinates admit exact closed forms. Since the $z$-values are a permutation of centered integers $\{1-\mu,\ldots,n^{2}-\mu\}$ where $\mu=(n^{2}+1)/2$, and the $x$- and $y$-coordinates form symmetric grids, standard calculations yield:
$\mathrm{Var}(Z)=\frac{n^{4}-1}{12},\qquad\mathrm{Var}(X)=\mathrm{Var}(Y)=\frac{n^{2}-1}{12}.$ (16)
We verified these formulas computationally for $n=3,4,5,7$ with exact agreement. Combined with the zero off-diagonal covariances from Proposition 3.6, this completely determines the covariance matrix for Magic Gems of any order.
Higher-Order Moment Vanishing.
The zero-covariance property $\mathbb{E}[XZ]=\mathbb{E}[YZ]=0$ extends to all higher powers. Since $X$ depends only on the column index $j$, we have
$\mathbb{E}[X^{k}Z]=\frac{1}{n^{2}}\sum_{j}X_{j}^{k}\cdot\left(\sum_{i}Z_{ij}\right)=\frac{1}{n^{2}}\sum_{j}X_{j}^{k}\cdot(C_{j}-n\mu),$
where $C_{j}$ is the sum of column $j$. For magic squares, $C_{j}=M(n)=n\mu$ for all $j$, so $\mathbb{E}[X^{k}Z]=0$ for any power $k$. Similarly, $\mathbb{E}[Y^{k}Z]=0$ follows from equal row sums. We verified $\mathbb{E}[X^{2}Z]=\mathbb{E}[Y^{2}Z]=\mathbb{E}[X^{3}Z]=\mathbb{E}[Y^{3}Z]=0$ for all tested magic squares across orders $n=3,4,5,7$.
Energy Landscape Local Minima.
Exhaustive enumeration for $n=3$ reveals the fine structure of the energy landscape. Under the complete energy $E_{\mathrm{full}}$, there are exactly 24 local minima: 8 global minima (the magic squares with $E_{\mathrm{full}}=0$) and 16 non-global local minima where no single swap decreases energy but $E_{\mathrm{full}}>0$. Under the simpler low-mode energy $E_{\mathrm{low}}$, the landscape is more rugged, with 344 local minima (8 global, 336 non-global). This difference reflects the fact that $E_{\mathrm{full}}$ has more terms ($2n$ vs. 4), providing more “escape directions” from non-magic configurations.
@@ 4.1.4 Extended Statistical Properties
4.2 The 4×4 Case and the Low-Mode Relaxation
@@
Proposition 3.6 makes no distinction between even and odd orders, predicting $\mathrm{Cov}(X,Z)=\mathrm{Cov}(Y,Z)=0$ for all magic squares. Our computations confirm this prediction for the $4\times 4$ case: all magic squares achieve exactly zero aggregate covariance to machine precision.
However, as discussed in Section 3.4.2, the four-term “low-mode” energy $E_{\mathrm{low}}$ does not characterize magic squares for $n\geq 4$. Explicit counterexamples can be constructed: a $4\times 4$ arrangement with all four low-mode covariances equal to zero (hence $E_{\mathrm{low}}=0$), yet with row and column sums far from the magic constant. Such arrangements achieve zero covariance with the aggregate position coordinates $X$ and $Y$ by balancing the deviations in a first-moment sense, without requiring equal sums.
To contrast the two energies, Figure 4 presents a scatter plot comparing $E_{\mathrm{low}}$ and $E_{\mathrm{full}}$ across 5,000 random $4\times 4$ arrangements plus magic squares. Magic squares (gold stars) sit at the origin under both measures. The scatter reveals that the two energies measure genuinely different aspects of arrangement structure: small $E_{\mathrm{low}}$ does not guarantee small $E_{\mathrm{full}}$.
Figure 4: Comparing low-mode and complete energies for $n=4$. Scatter plot of $E_{\mathrm{low}}$ (4-term aggregate) versus $E_{\mathrm{full}}$ (using $2n$ individual indicators) for 5,000 random arrangements (blue dots) and magic squares (gold stars at origin). The inset zooms near the origin. The two energies are correlated but measure different balance properties; only $E_{\mathrm{full}}=0$ characterizes magic squares for $n\geq 4$.
4.3 Geometric Structure of the $n=4$ Magic Square Population
Beyond validating the zero-covariance theorem, the $n=4$ case provides an opportunity to explore geometric diversity within the magic square population. The 880 essentially different $4\times 4$ magic squares, while all satisfying the same covariance conditions, exhibit varying polyhedral structures and perturbation resistance.
The comprehensive perturbation analysis (Figure 2) revealed a striking trimodal distribution of minimum gaps for $n=4$, suggesting that the 880 squares cluster into geometric subclasses with different stability properties. This geometric diversity, combined with the trimodal gap structure, indicates that the 880 essentially different $4\times 4$ magic squares may admit a meaningful geometric classification beyond $D_{4}$ equivalence.
@@ 4.4 Cross-Order Comparison and Scaling Analysis
The zero-covariance theorem holds universally across orders, but the geometric and energy landscape properties evolve systematically with $n$. Figure 5 presents a unified view of Magic Gems across orders $3$, 44, and $5$, along with key scaling metrics.
Figure 5: Cross-order comparison and scaling analysis. Top row (a–c): Side-by-side comparison of $3\times 3$, $4\times 4$, and $5\times 5$ Magic Gems. The odd-even-odd progression reveals structural differences (e.g., central vs. off-center grids) while the zero-covariance property persists across all orders. Bottom row (d–f): Scaling properties. (d) Convex hull volume increases with order, reflecting the larger range of $z$-values. (e) The fraction of points on the hull boundary decreases as interior points proliferate. (f) The magic constant $M(n)=n(n^{2}+1)/2$ grows quadratically.
Several systematic patterns emerge. The convex hull volume grows rapidly with $n$, from 18.0 for $n=3$ to 276.7 for $n=5$, reflecting both the expanding $z$-value range (4-4 to $4$ for $n=3$, versus 12-12 to $12$ for $n=5$) and the increasing number of vertices. The fraction of points lying on the hull boundary decreases from $8/9\approx 0.89$ for $n=3$ to roughly $0.64$ for $n=5$, as the hull interior becomes more populated.
Table 2 summarizes key properties across orders, demonstrating both the growth in complexity and the invariance of the zero-covariance characterization.
@@
Hull vertices (typical) 8 12 16
@@
Table 2: Comparison of Magic Gem properties across orders 3, 4, and 5. All orders exhibit the zero aggregate covariance property (Proposition 3.6). For $n=3$, the low-mode energy $E_{\mathrm{low}}$ suffices to characterize magic squares; for $n\geq 4$, the complete energy $E_{\mathrm{full}}$ is required (Remark 3.10).
The even-order case ($n=4$) demonstrates that the zero-covariance property does not depend on having a central cell at the origin. The algebraic proof—that equal row and column sums imply vanishing aggregate covariances—holds regardless of the grid’s parity, confirming the universality of the framework. However, the insufficiency of $E_{\mathrm{low}}$ for $n\geq 4$ (Figure 4) underscores the necessity of the complete indicator-based energy $E_{\mathrm{full}}$ for a rigorous characterization.
@@ 4.5 Large-Scale Energy Landscape Analysis
With the Complete Energy Characterization Theorem (Theorem 3.17) established, we turn to exploring the structure of the energy landscape and its scaling properties. Leveraging modern multi-core hardware (Mac Studio M5 with 128GB RAM), we analyzed over 460 million arrangements across orders $n=3,4,5$: exhaustive enumeration for $n=3$ (362,880 arrangements), 60 million samples for $n=4$ (from 20.9 trillion total), and 400 million samples for $n=5$ (from approximately $15.5\times 10^{24}$ total).
Note that this sampling uses the low-mode energy $E_{\mathrm{low}}$ (Definition 3.18), which coincides with the complete energy for $n=3$ but defines a relaxed notion for larger orders. The analysis characterizes the landscape structure and the rarity of near-zero configurations.
@@ 4.5.1 Energy Distribution Evolution
Figure 6 presents a comprehensive 9-panel analysis of the energy landscape across orders $n=3,4,5$, revealing a dramatic evolution in landscape structure. For $n=3$, the exhaustive enumeration confirms that exactly eight arrangements (the Lo Shu and its $D_{4}$ variants) achieve zero energy, with a clear gap separating them from all other arrangements.
Figure 6: Low-mode energy ($E_{\mathrm{low}}$) landscape analysis across orders $n=3,4,5$. Based on exhaustive enumeration ($n=3$: 362,880), and extensive sampling ($n=4$: $6\times 10^{7}$; $n=5$: $4\times 10^{8}$). Top row: Individual distributions (log scale) showing modal energy shifting from 3.12 to 15.12. Middle row: Overlaid comparison, normalized PDFs, and cumulative distributions. Bottom row: Mean energy with error bars, landscape entropy (varying modestly across orders, $H\approx 4.3$$5.0$), and peak shift with quadratic fit ($R^{2}>0.99$). The red dashed line marks $E_{\mathrm{low}}=0$. For $n=3$, this coincides with magic squares; for $n\geq 4$, it also includes non-magic “low-mode balanced” configurations (see Remark 3.10). The complete energy $E_{\mathrm{full}}$ characterizes magic squares exactly for all $n$.
The key statistics from this analysis are summarized in Table 3. For $n=3$, exhaustive enumeration confirms that exactly 8 of 362,880 arrangements achieve zero low-mode energy, corresponding precisely to the $D_{4}$ variants of the Lo Shu square. For $n=4$ and $n=5$, no zero-energy configurations were found by random sampling—however, this should not be interpreted as evidence that magic squares are the only zeros of $E_{\mathrm{low}}$. As discussed in Remark 3.10, non-magic zeros exist for $n\geq 4$, but they are sufficiently rare that uniform random sampling is unlikely to discover them.
@@
Zero-energy fraction $2.2\times 10^{-5}$ $10^{-7} <$$10^{-8}$
Low-mode zeros found 8 0 0
Table 3: Low-mode energy ($E_{\mathrm{low}}$) landscape statistics from large-scale computational analysis. Based on exhaustive enumeration for $n=3$ and extensive random sampling for $n=4,5$. For $n\geq 4$, the low-mode energy has non-magic zeros (see Remark 3.10), but these are too rare for random sampling to detect. The complete energy $E_{\mathrm{full}}$ correctly characterizes magic squares for all orders.
@@ 4.5.2 Scaling Laws
where $p_{k}$ is the fraction of arrangements falling in energy bin $k$ (using 200 equally-spaced bins covering the observed energy range). Remarkably, the landscape entropy remains within a narrow band across orders ($\approx 4.3$$5.0$), decreasing from $n=3$ to $n=4$ and then stabilizing. This suggests that the relative structure of the landscape is scale-invariant: the difficulty of navigating the landscape by local search methods does not fundamentally change with order, though the absolute energy scales increase substantially.
@@ 4.5.3 Scaling Properties Summary
Figure 7 provides a compact four-panel summary of how energy landscape properties scale with magic square order, synthesizing the key quantitative findings. The energy landscape “spreads” as $n$ increases, while the zero-energy position remains the isolated location of magic squares across all orders.
Figure 7: Low-Mode Energy Scaling Properties. (A) Peak low-mode energy follows quadratic scaling with $R^{2}>0.99$. (B) Mean energy with standard deviation error bars, showing increasing variance. (C) Landscape entropy varies modestly across orders ($H\approx 4.3$$5.0$), suggesting scale-invariant structure despite dramatic energy range expansion. (D) Sample sizes tested (blue, log scale) versus magic square probability (red, log scale), demonstrating the computational challenge of finding magic squares at larger $n$. Note: These statistics characterize the low-mode energy $E_{\mathrm{low}}$ landscape; the complete energy $E_{\mathrm{full}}$ has qualitatively similar scaling behavior.
@@ 4.5.4 Computational Verification Summary
This large-scale analysis provides computational insight into the energy landscape structure:
@@
    Complete Characterization Theorem: Proven analytically in Theorem 3.17—for all $n\geq 3$, $E_{\mathrm{full}}(S)=0$ if and only if $S$ is a magic square. This follows directly from the row/column covariance characterization (Proposition 3.15, Proposition 3.12).
@@
    Low-Mode Characterization (n=3 only): Exhaustively verified—exactly 8 of 362,880 arrangements have zero low-mode energy $E_{\mathrm{low}}$, all of which are magic squares. Notably, 760 arrangements satisfy $\mathrm{Cov}(X,Z)=\mathrm{Cov}(Y,Z)=0$, but only the 8 magic squares additionally satisfy the diagonal conditions.
@@
[…]
@@
    Scaling Behavior: The low-mode energy landscape exhibits systematic scaling, with the peak energy appearing to grow approximately quadratically with order based on the three data points available ($n=3,4,5$).
  5. 5.
@@
These computational results complement the theoretical analysis of Section 3, characterizing the structure of the energy landscape and establishing quantitative scaling laws for future investigation.
@@ 5.1 The Meaning of Zero Covariance
The zero-covariance property of magic squares (Proposition 3.6) admits several complementary interpretations that illuminate their structure.